OpenStudy (anonymous):

A bookcase is 3 feet high and 3 feet wide. Two braces are going to be built to diagonally cross the back of the case. How long is the piece of wood that is needed to build each brace ? A) 3 3 feet B) 3 2 feet C) 3 feet D) 6 feet

OpenStudy (anonymous):

42

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

Use the pythagorean theorem to solve for x a^2 + b^2 = c^2 3^2 + 3^2 = x^2 ... I'll let you finish

OpenStudy (anonymous):

thanks jim

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

thanks jim